How many liters of a 50% alcohol solution must be mixed with 90 liters of a 90% solution to get a 80% solution? Answer is 30L and I keep getting REALLY high numbers.
Answer by mananth(16946)

------ percent ---------------- quantity
Alcohol I 90----------------90 liters
Alcohol II 50----------------x liters
Total 80---------------90+x liters

90*90+50 *x=80(90+x)
8100+50x =7200+80x
50x-80x= 7200-8100
-30x=-900
/ -30
x=30 liters 50 % alcohol II